OpenStudy (helpmeee123):

a nar contains quarters and dimes. There are 15 more quarters than dimes. The total value of the coin is $23. Which system of equations can be used to find the number of the quarters q and the number of dimes d. A.) q=d-15 0.25 + 0.1d= 0.23 B.) q = d+15 0.25q + 0.1d = 0.23 C.) d=q +15 0.25q +0.1d= 23 D.) q=d +15 0.25q + 0.1d= 23 Please show me a work. Thank you thank uoi thank you

OpenStudy (anonymous):

Assume q = d but because there are 15 more quarter than dimes so it has to add 15 into d to make it equal to q

OpenStudy (anonymous):

there fore q = d + 15

OpenStudy (anonymous):

\[\{10 d+25 q=2300,q-15=d\} \] d=55, q=70

OpenStudy (anonymous):

the total value for each kind calculate by multiply the number coin with its value 0.25q + 0.1d = 23 Now you simplify it to get the second equation. Now you have it

OpenStudy (anonymous):

D is the answer

OpenStudy (anonymous):

Anyone else calculate d=55 and q=70 using dollars?